a. The manufacturer of aspirin claims that the proportion
of headache sufferers who get relief with
just two aspirins is 53%. What is the probability
that in a random sample of 400 headache sufferers,
less than 50% obtain relief? If 50% of the
sample actually obtained relief, what does this
suggest about the manufacturer’s claim?
b. Repeat part (a) using a sample of 1,000
1
Expert's answer
2020-12-10T19:45:30-0500

From the claim we obtain that the probability to get a relief for a person is "0.53".

a).The task is to find the probability that less than 200 sufferers will get relief. We will use the binomial distrubution with parameters 400 and "0.53" (see https://mathworld.wolfram.com/BinomialDistribution.html). The probability that "k" sufferers will get relief is: "p_k=P(X=k)=\frac{400!}{k!(400-k)!}(0.53)^k(0.47)^{400-k}"

"X" denotes a random variable with the binomial distribution.

The aim is to find the probability: "\sum_{k=0}^{199}p_k=\sum_{k=0}^{199}\frac{400!}{k!(400-k)!}(0.53)^k(0.47)^{400-k}\approx"

"\approx0.1053"

The latter was rounded to 4 decimal places. We used the following command in Maple 13 to compute the sum:

sum(400!/(k!*(400-k)!)*(0.53)^k*(0.47)^(400-k),k=0..199);


Suppose that 50% obtained relief. In case we allow for the value to differ on "3\%" from "53\%" we can say that the claim is true. However, if we require the difference to be less than "3\%", then the claim is false.

b). The task is to find the probability that less than "500" sufferers will get relief. We use again the binomial distribution with parameters 1000 and "0.53". The aim is to compute

"\sum_{k=0}^{499}\frac{1000!}{k!(1000-k)!}(0.53)^k(0.47)^{1000-k}\approx0.0267"


We computed it via the following command in Maple 13:

sum(1000!/(k!*(1000-k)!)*(0.53)^k*(0.47)^(1000-k),k=0..499);


The latter is rounded to 4 decimal places.

Answer: a). 0.1053; b). 0.0267